OpenStudy (anonymous):

What is 7/20 as a decimal

OpenStudy (anonymous):

Thats easy how old r u?

OpenStudy (anonymous):

Divide 7 by 20

OpenStudy (anonymous):

i have an idea, write \[\frac{7}{20}=\frac{7\times 5}{20\times 5}=\frac{35}{100}\] and then the answer should be more clear

OpenStudy (anonymous):

It's 0.35